OpenStudy (anonymous):

How do I simplify: y-1=2/3(x-4) The answer is y=2/3x-5/3 I just don't know how to get there. Thanks

OpenStudy (anonymous):

y-1+1= 2/3(x-4)+1 y=2/3(x-4)+1 y=2/3x -2/3 -3 y=2/3x-5/3

OpenStudy (anonymous):

How does the 2/3 distribute into the (x-4)? I can't seem to do it

OpenStudy (anonymous):

since the x -4 is in parenthesis, u distribute 2/3 into x and also 2/3 into -4

OpenStudy (anonymous):

2/3 into -4 is -2.66666

OpenStudy (anonymous):

you are just simplifying the problem,,, not trying to figure it out.

OpenStudy (anonymous):

I see I guess I dont' get how to simplify it.

OpenStudy (anonymous):

y-1= 2/3(x-4) y=1 + 2/3(x) - 2/3(4) y=1+2x/3 -8/3 y=2x/3 +1 - 8/3 y= 2x/3 + (3-8)/3 y=2x/3 - 5/3 got that!!!!!!!!!!

OpenStudy (anonymous):

got it! Thanks

OpenStudy (phi):

when distributing a fraction you can do it two ways: use decimals or use fractions if you use decimals 2/3 = 0.6666... the 6's go on forever (people put a line over the 6 to show this) and 0.6666(x-4)= 0.6666*x - 2.6666... but for these problems it looks nicer to leave everything as a fraction when you multiply fractions, you multiply top times top and bottom by bottom if you multiply a fraction times a whole number, assume the whole number is over 1 (its bottom is 1) \[ \frac{2}{3}(x-4) = \frac{2}{3}\cdot\frac{x}{1} + \frac{2}{3}\cdot \frac{-4}{1} \] \[ \frac{2x}{3} - \frac{8}{3} \]